A balloon is rising vertically above a level, a straight road a constant rate 1 m/s. Just when the balloon is 65m above the ground, a bicycle moving at a constant rate of 17 m/s passes under it. How fast is the distance b/w the bicycle and balloon increasing 3 seconds later?


A)11 m/s


B) 9 m/s


C) 13 m/s


D) 15 m/s

Grade:

1 Answers

Arun Kumar IIT Delhi
9 years ago
we do this by coordinate geometry baloon (0,65+1*t)
bicycle (17t,0)
z^2=289*t^2+(65+t)^2
2zdz/dt=289*2*t+2(65+t)
we can dz/dt=11